40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15714     Entry time: Mon Dec 7 14:32:02 2020
Author: gautam 
Type: Update 
Category: LSC 
Subject: New demod phases for POX/POY locking 

In favor of keeping the same servo gains, I tuned the digital demod phases for the POX and POY photodiode signals to put as much of the PDH error signal in the _I quadrature as possible. The changes are summarized below:

POX / POY demod phases
PD Old demod phase [deg] New demod phase [deg]
POX11 79.5 -75.5
POY11 -106.0 116.0

The old locking settings seem to work fine again. This setting isn't set by the ifoconfigure scripts when they do the burt restore - do we want it to be?

Attachments #1 and #2 show some spectra and TFs for the POX/POY loops. In Attachment #2, the reference traces are from the past, while the live traces are from today. In fact, to have the same UGF as the reference traces (from ~1 year ago), I had to also raise the digital servo loop gain by ~20%. Not sure if this can be put down to a lower modulation depth - at least, at the output on the freq ref box, I measured the same output power (at the 0dB variable attenuator gain setting we nominally run in) before and after the changes. But I haven't done an optical measurement of the modulation depth yet. There is also a hint of lesser phase available at ~100 Hz now compared to a year ago.

Attachment 1: POX_POY_OLTF.pdf  42 kB  Uploaded Mon Dec 7 16:03:40 2020  | Hide | Hide all
POX_POY_OLTF.pdf
Attachment 2: POX_POY_spectra.pdf  125 kB  Uploaded Mon Dec 7 16:04:22 2020  | Hide | Hide all
POX_POY_spectra.pdf
ELOG V3.1.3-